OpenStudy (anonymous):

An object is dropped from a height of 25 meters. At what velocity will it hit the ground? a 7.0 m/s b 11 m/s c 22 m/s d 49 m/s e 70 m/s

OpenStudy (anonymous):

c. 22m/s \[\sqrt{2\times9.8\times25}\]

OpenStudy (astrophysics):

To find the final velocity for a falling object, we actually get it from total energy for a falling object, \[v_f = \sqrt{2gh}\]

OpenStudy (shamim):

We know the formula v^2=u^2+2*g*h

OpenStudy (shamim):

v=?

OpenStudy (shamim):

Given u=0 g=9.8m/s h=25m